OpenStudy (anonymous):

There are a total of 255 dogs, cats, fish in a pet store. The ratio of dogs to cats is 3:5 and ratio of cats to fish is 2:7. How many fish are in the pet store?

OpenStudy (kropot72):

Let the number of cats be c, the number of dogs be d and the number of fish be f. \[\frac{d}{c}=\frac{3}{5}\] Therefore \[d=\frac{3}{5}c\ ........(1)\] Also \[\frac{c}{f}=\frac{2}{7}\] Therefore \[f=\frac{7}{2}c\ ...........(2)\] Equations (1) and (2) give the numbers of dogs and fish in terms of the number of cats. \[c+d+f=255\ ...........(3)\] Plugging in the expressions for d and f from equations (1) and (2) into equation (3) we get \[c+\frac{3}{5}c+\frac{7}{2}c=255\ ........(4)\] Now you can use equation (4) to find the number of cats, and then by substitution into equation (2) the number of fish can be found.

OpenStudy (anonymous):

Thank you so much~

OpenStudy (kropot72):

You're welcome :)